# Knowledge-Driven Feature Selection and Engineering for Genotype Data with Large Language Models

### v1

@@ Methods
In Table 1, we examine the performance of different LLMs in our framework, repeating the experiment five times with cross-validation. We find that other models, including the relatively weaker GPT-3.5-turbo, are highly competitive. This may be surprising but this further suggests that there is a limited effectiveness to the types of interaction terms we enable, which ranges from conditional rules to multiplicative expressions, and this may mean that these interactions are within the reasoning capabilities of the models we study. These results also align with the models’ shared awareness of the variants and their relevance to the task (see Figure 5); it may mean that knowledge demonstration is enough for the LLM to provide utility as a weak learner.
Table 1: FreeForm using different models for featuring engineering on Hearing Loss (16-shot) and Ancestry (20-shot). LR: Logistic Regression. RF: Random Forest.
Hearing Loss Ancestry
Model GPT-3.5-turbo Llama-3.1-405B GPT-4o GPT-3.5-turbo Llama-3.1-405B GPT-4o
LR AUC (Std) 0.5064 (0.10) 0.5113 (0.11) 0.5145 (0.11) 0.9546 (0.01) 0.9562 (0.02) 0.9572 (0.01)
RF AUC (Std) 0.5402 (0.11) 0.5444 (0.11) 0.5438 (0.11) 0.9514 (0.01) 0.9520 (0.01) 0.9527 (0.01)
@@
In our main results of FreeForm, we use a simple serialization strategy like the following: “$s_{1}$ is $x^{i}_{1}$. $s_{2}$ is $x^{i}_{2}$… Answer: $y_{i}$”. In this ablation study, we find that using a more elaborate schema like “The $s_{1}$ variant of the person has $x^{i}_{1}$ minor alleles ...” does not make a difference. While this may be surprising given the existing efforts on exploring serialization strategies 46, it could be that foundational models are robust to formatting and that these strategies are less effective when used with in-context learning. Furthermore, our findings show that the augmentation of gene information is not significant. Our experimentation is limited to providing the genes associated with the variant. This may be redundant information for the LLM but other strategies are not straightforward; augmenting literature for a variant is challenging due to the lack of relevance it may have to the task. This will be an important avenue to explore for future work.
@@
The genomic ancestry dataset we used is acurated version of the full dataset, after applying rigorous quality control. While this ensures data reliability, it can lead to the omission of significant genetic variants. To address this concern, we ask GPT-4o to suggest fifteen SNPs, and we add these onto the curated dataset—SNPs such as rs671 or rs2814778 with clear causal relationships or statuses as standard AIMs were suggested. Our LLM-driven selection methods were able to recover most of these variants while the data-driven methods could not, and we found that these variants contributed largely to the performance gap we observed for feature selection in the genomic ancestry dataset. We find it concerning that data-driven methods failed to identify many of these variants despite their predictive power; training a Logistic Regression model with the GPT-4o-suggested SNPs achieved an average of 0.94 AUC on 20 shots whereas a set of fifteen SNPs selected by PCA on the original dataset achieved 0.78 AUC. LLMs could be a promising method to mitigate the issues of quality control, a frequent issue in this domain, providing a potentially more robust, automatable method than manually imputing well-known variants from the literature.
Figure 5: Comparison between different LLM models on their knowledge of the SNP rs671 relating to genomic ancestry. Red text indicates a hallucination, which was only observed in the case of the Llama 2 7B model. Every other model correctly identifies the genetic association.
@@ Discussion
We present FreeForm which advances the state-of-the-art in LLM-based, few-shot tabular learning and we novelly apply our LLM-driven framework to genotype data. FreeForm goes beyond the typical usage of LLMs for inference and aids the process of feature selection and engineering, tackling the issues of high dimensionality, limited data samples, and interpretability. In our exploration of genotype data, we find that LLMs have a robust knowledge of genetic variants, demonstrating state-of-the-art performance across datasets with different variant ID schemas.
Our framework notably has a few key advantages over existing approaches besides performance: (1) its model-agnostic nature and (2) FreeForm requires no inference costs as the features need to be engineered only once at training, compared to LLM-only methods that require large computational costs at inference time. Moreover, the entire pipeline can be conducted with just API access to the model, costing around a dollar to run the entire pipeline. Currently, top-performing LLMs either only offer access through inference APIs or require significant resources to load or fine-tune.
@@
However, FreeForm still has room for growth; our study suggests we allow a naive range of interaction types in feature engineering, our scalable approach to feature engineering can integrate more advanced strategies, and our retrieval augmentation is limited to gene information. Our future goals are to expand its capabilities in generating novel features, improve the extraction of knowledge intrinsic to LLMs, better augment task-specific knowledge from APIs such as PubMed, and integrate interpretability or explainability with the assistance of LLMs to enable functionality such as causal discovery. Recent studies have demonstrated the potential for LLMs in the discovery of new gene sets or causal genes 25; 26, where LLMs effectively interpolate across the vast corpus of scientific literature they are trained on. Furthermore, we see promising future work in developing our approaches in feature selection and nomination, where we used the LLM to suggest predictive features. The controllability of LLMs through prompting opens up interesting possibilities, such as the nomination of features that better represent diverse populations, thereby mitigating biases that data-driven methods can exacerbate. A limitation of our study is that several assumptions were made in the evaluation, such as the selection of fifteen features, to showcase our pipeline. Thus, we plan to expand our evaluation to more scenarios and more phenotypes, such as Alzheimer’s disease, to demonstrate its robust utility.

### v2

@@ Results
In Table 1, we examine the performance of various LLMs in our framework, repeating the experiment five times with cross-validation. For feature selection, gpt-4o performs better as expected, except for hearing loss, further suggesting limited knowledge of the relevant variants across the LLMs. It remains unclear whether this advantage arises from the LLMs’ ability to identify relevant features or their depth of knowledge (e.g. while an LLM may recognize a variant, its expertise regarding that variant may vary). For feature engineering, the weaker models are highly competitive. This is surprising; this may indicate that the interaction terms we enable (e.g., conditional rules, multiplicative expressions) are within the reasoning capabilities of all studied models. Alternatively, it could reflect all the models’ inability to capture epistasis. Either way, these results align with the shared awareness of variants among models (see Figure 5), suggesting that even basic knowledge can enable LLMs to act as weak learners.
Table 1: FreeForm using different models. For feature selection, hierarchical selection is used across all models. We limit the analysis to 16-shots (Hearing Loss) and 20-shots (Ancestry). LR: Logistic Regression. RF: Random Forest. Values shown are AUROC (AUC). Standard deviation is in parenthesis.
Model Feature Selection Feature Engineering
Hearing Ancestry Hearing Ancestry
LR RF LR RF LR RF LR RF
GPT-3.5-turbo 0.524 (0.03) 0.539 (0.04) 0.785 (0.03) 0.765 (0.04) 0.506 (0.10) 0.540 (0.11) 0.955 (0.01) 0.951 (0.01)
Llama-3.1-405B 0.511 (0.11) 0.544 (0.11) 0.956 (0.02) 0.952 (0.01)
GPT-4o 0.506 (0.09) 0.544 (0.10) 0.943 (0.01) 0.943 (0.02) 0.514 (0.11) 0.543 (0.11) 0.957 (0.01) 0.953 (0.01)
@@
In our main results of FreeForm, we use a simple serialization strategy like the following: “$s_{1}$ is $x^{i}_{1}$. $s_{2}$ is $x^{i}_{2}$… Answer: $y_{i}$”. In this ablation study, we find that using a more elaborate schema like “The $s_{1}$ variant of the person has $x^{i}_{1}$ minor alleles ...” does not make a difference. While this may be surprising given the existing efforts on exploring serialization strategies 42, large foundation models may be robust to such formatting. Furthermore, our findings show that the augmentation of gene information is not significant. Our experimentation is limited to providing the genes associated with the variant. This may be redundant information for the LLM but other strategies are not straightforward; augmenting literature for a variant is challenging due to the lack of relevance it may have to the task. This will be an important avenue to explore for future work.
@@
The genomic ancestry dataset we used is a curated version of the full dataset, after applying rigorous quality control. While this ensures data reliability, it can lead to the omission of significant genetic variants. To address this concern, we ask GPT-4o to suggest fifteen SNPs; SNPs such as rs671 or rs2814778 with clear causal relationships or statuses as standard AIMs were usually suggested. These SNPs originally existed in the database but were omitted in the curation of the 10K version so we inserted them back. Our LLM-driven selection methods were able to recover most of these variants while the data-driven methods could not, and we found that these variants contributed largely to the performance gap we observed for feature selection in the genomic ancestry dataset. We find it concerning that data-driven methods failed to identify many of these variants despite their predictive power. In additional analysis, training a Logistic Regression model with the GPT-4o-suggested SNPs alone achieved an average of 0.94 AUC on 20 shots whereas a set of fifteen SNPs selected by PCA on the original dataset achieved 0.78 AUC. LLMs could be a promising method to mitigate the issues of quality control, a frequent issue in this domain, providing a potentially more robust, automatable method than manually imputing well-known variants from the literature.
Figure 5: Comparison between different LLM models on their knowledge of the SNP rs671 relating to genomic ancestry. Red text indicates a hallucination, which was only observed in the case of the Llama 2 7B model.
@@ Discussion
We present FreeForm which advances the state-of-the-art in LLM-based, few-shot tabular learning and we novelly apply our LLM-driven framework to genotype data. FreeForm goes beyond the typical usage of LLMs for inference and aids the process of feature selection and engineering, tackling the issues of high dimensionality, limited data samples, and interpretability. Furthermore, we find that LLMs have a robust knowledge of genetic variants, demonstrating state-of-the-art performance across different variant ID schemas, showcasing the promise of LLMs in genetics.
Our framework notably has several key advantages over existing approaches besides performance: (1) it is model-agnostic, (2) it scales well to higher dimensions, and (2) it incurs no inference costs, as features are engineered once during training, unlike LLM-only methods with high computational costs at inference time. Moreover, the entire pipeline can be executed using API access, costing approximately one dollar to run, with the majority of the pipeline completing within minutes, aside from the initial filtering step in feature selection.
@@
However, FreeForm has room for growth. The current results suggest we either allow a naive range of variants interaction types in feature engineering, or none of the models are capable of capturing epistasis. Our retrieval augmentation is also limited to gene information. Additionally, the LLM’s input could be enhanced by including high-level feature statistics typically considered in this domain. Future work remains in expanding its capabilities to generate novel features, improve the extraction of knowledge intrinsic to LLMs, better augment task-specific knowledge from APIs such as PubMed, and integrate further interpretability or explainability. Recent studies have demonstrated the potential for LLMs in the discovery of new gene sets or causal genes 24; 25, where LLMs effectively interpolate across the vast corpus of scientific literature they are trained on. We find this approach promising for addressing the issue of multicollinearity, which we acknowledged but did not resolve in this work. Furthermore, we see promising future work in developing feature nomination, where we used the LLM to suggest predictive features. The controllability of LLMs through prompting opens up interesting possibilities, such as the nomination of features that better represent diverse populations, thereby mitigating biases that data-driven methods can exacerbate. One limitation of our study is that a few assumptions were made in the evaluation, including the fixed selection of fifteen features, to showcase our pipeline. Thus, we plan to expand our evaluation to more scenarios and more phenotypes, such as Alzheimer’s disease, to demonstrate its robust utility.
As LLMs advance in domain expertise, potentially surpassing humans 43, their potential to revolutionize bioinformatics becomes increasingly imminent. While our study demonstrates LLMs’ excellence in low-shot regimes, we acknowledge such scenarios are rare in practice. We anticipate, however, these capabilities will scale as foundational models advance and domain-specific LLMs develop. Recent efforts, such as fine-tuning foundational models on literature 43 or augmenting LLMs with knowledge graphs 44, are making progress towards this. While our framework focuses on feature selection and engineering, our work serves as a prototype, showcasing the potential of LLMs in genetics.